Solve the inequality
100-x^2>=0
Answer choices:
A.) x<=-50 or x>= 50
B.) x<=10
C.) x<=-10 or x>=10
D.) -10<=x<=10
Solve the inequality
Solve the inequality
100-x^2>=0
Answer choices:
A.) x<=-50 or x>= 50
B.) x<=10
C.) x<=-10 or x>=10
D.) -10<=x<=10

100-x^2>=0 SET TO ZERO
(10+X)(10-X) >= 0 EXPAND EQUATIONS
10-X>=0 OR 10+X >=0 MAKE SURE TO PLUG IN AND DRAW THE LINE TO VARIFY
X<=10 OR X>=-10 -10<=X<=+10 ANSW IS D

Answer by chiexpert(48)

Solve the inequality
Solve the inequality
100-x^2>=0
Answer choices:
A.) x<=-50 or x>= 50
B.) x<=10
C.) x<=-10 or x>=10
D.) -10<=x<=10

100-x^2>=0 SET TO ZERO
(10+X)(10-X) >= 0 EXPAND EQUATIONS
10-X>=0 OR 10+X >=0 MAKE SURE TO PLUG IN AND DRAW THE LINE TO VARIFY
X<=10 OR X>=-10 -10<=X<=+10 ANSW IS D